# Rapid Rotation of Polarization Orientations in PSR B1919$+$21’s Single Pulses: Implications On Pulsar’s Magnetospheric Dynamics

### v1

@@ 3.4 Rapid rotating polarization orientations in single pulses
Next we directly investigate the individual pulses. Several pulses are presented in Figure 4, with polarization evolution tracks on Poincaré spheres plotted alongside. The polarization states of individual pulses still seem irregular, with PA tracks appearing more arbitrary. Not all pulses show “discontinuities” around all three vertical grey dashed lines. Orthogonal mode jumps appear in some pulses’ central longitude ranges (e.g., #274 in Figure 4, at longitude $\sim 5^{\circ}$). It’s worth noticing that some pulses present monotonically rotating long PA curves in some longitude ranges (see #7, #275, #555 and #2565 in Figure 4 for example). The PA in those pulses rotates by over $180^{\circ}$ or even $360^{\circ}$ across longitudes, accompanied with manifest oscillations of circular polarization (or ellipticity angle) values around 00. The polarization states at these pulse longitudes seem to follow long circular tracks on Poincaré spheres.
Figure 4: (i) Upper 6 plots: 6 single pulses with their polarization states shown on Poincaré spheres. In “Intensity” panels, the lines’ meaning are same as Figure 1. “$P/I$” panels show the polarization fraction $\sqrt{Q^{2}+U^{2}+V^{2}}/I$ with errorbars on each longitudes (only longitudes where $\sigma_{\mathrm{PA}}\leq 5$ are chosen, same for “$\psi$” panels and “$\chi$” panels). “$\psi$” panels show PAs with errorbars, and “$\chi$” panels show EAs with errorbars, on each chosen longitudes. PA and EA dots are also plotted on Poincaré spheres under Hammer-Aitoff projection. The change of color from darker to lighter of PA and EA dots represents the increase in longitude. (ii) Lower 3 plots: 3 single pulses with PA, EA and total intensity $I$ plotted in two frequency subbands (1000-1125 MHz in green and 1375-1500 MHz in orange). In the $I$ panel, the intensity curves are normalized to the maximum values separately for two subbands.
Such rapid rotating polarization patterns are not rare for B1919$+$21, with over 1/3 pulses showing quasi-monotonic rapid rotations of PA in some longitude ranges. This phenomenon seems to depend only very weakly on frequency (see Figure 4 for three examples), at least in our observational bandwidth. In addition, considering PA curves as functions of pulse longitudes, most quasi-monotonic PA curves in single pulses have a generally negative slope (e.g., #7, #555 and #2565 in Figure 4), while only very few of them have a positive slope (e.g., #275 in Figure 4). This asymmetry in PA curve slopes’ distribution might have some physical implications, which we will discuss later. Finally, we’d like to point out that the PA curves between two adjacent pulses (e.g., #274 and #275 in Figure 4) can change vastly.
@@ 4 Attempts at modeling
Among the observed phenomena, the rapid rotating polarization orientations in single pulses are outstanding. The PA curves are far from the RVM types. OPM jumps in integrated profile and single pulses lead us to seek solutions from OPM related mechanisms. Coherent or partially coherent summation of orthogonal polarized waves (e.g., Dyks, 2019; Oswald et al., 2023), allowing only varying amplitude ratio, requires that the emitted radio wave’s polarization orientation should be initially rotating at different longitudes (to result in over $360^{\circ}$ rotation of PA), before propagating. The origin of such a large intrinsic rotation mechanism seems unclear. In this section, we seek solutions based on propagational effects in the magnetosphere. Accompanied by the PA curve, the circular polarization oscillation could indicate the change of phase lags between the wave modes across longitudes, which should be taken into consideration.
@@ 4.1 A phenomenological model: phase lag driven
To reproduce the observed polarization variations in single pulses, here we put forward a phenomenological model based on propagational effects. Based on theories like Cheng & Ruderman (1979); Lyubarskii & Petrova (1998); Petrova & Lyubarskii (2000), a pulsar magnetosphere could be divided into the adiabatic walking region (the magnetic field is strong enough and normal modes are linearly polarized) and the polarization limiting region (the weakening or deflection of local magnetic fields introduces more circular polarization components). Here we simply assume that two initially mutually orthogonal radio waves (O wave and E wave) propagate through two plasma layers. Directions of magnetic fields in two layers are fixed separately: in the first layer, the magnetic field is aligned with the O wave electric field vector; in the second layer, the magnetic field has an angle $\theta_{0}$ deflected from the first layer’s magnetic field. The phase lag between O wave and E wave increases by a value $\eta$ through propagating in the first layer. In the second layer, the original O and E waves are decomposed into new normal wave modes, between which the phase lag increases by $\delta$ through propagating. To include depolarization, we also follow Oswald et al. (2023) to introduce coherence $C$ into our modeling, and then a part of waves are added incoherently. The mathematical details of our model construction are in Appendix A. Schematic diagrams of the model described above are shown in Figure 5 (i).
@@
Figure 5: (i) A schematic diagram for the phenomenological model. The grey arrows represent magnetic fields. The green long arrow with label “$\boldsymbol{k}$” means the direction of wave propagation, where refraction is neglected. The blue rectangle with label “O” and the red parallelogram with label “E” represent O mode wave and E mode wave. (ii) A cross-sectional view of (i), and also a schematic diagram for the difference between two plasma layers. (iii) A pulse simulated from the phenomenological model. The horizontal axis is the number of sample bins, which is equivalent to the longitude. The red dash line, the blue line and the cyan line are linear, circular and total polarization fraction on each bin. The green and magenta dots are PA and EA on each bin. (iv) linear, circular, total polarization fraction, as well as PA and EA on each bin of pulse #555 in Figure 4.
There are 5 parameters in the model: (1) initial amplitude ratio $E_{\mathrm{O}}/E_{\mathrm{E}}$; (2) first layer’s phase lag $\eta$; (3) magnetic field deflection $\theta_{0}$; (4) second layer’s phase lag $\delta$; (5) coherence $C$. An example of a pulse simulated from the model is shown in Figure 5 (iii). For this pulse, the parameters are chosen as follows: (1) $E_{\mathrm{O}}/E_{\mathrm{E}}=1$; (2) $\eta$ ranges from 0 to $6\pi$ on each bin, from left to right; (3) $\theta_{0}=20^{\circ}$; (4) $\delta=\pi/4$; (5) $C=0.5$. Comparing with pulses in real data (e.g. #555 in Figure 5 (iv)), our modeling works well for understanding the polarization variations within individual pulses. The PA curve of the simulated pulse varies by $3\pi$ across the range monotonically, and the oscillating circular polarization (or EA) is also reproduced. For normal modes in pulsar radiation, when the phase lag between them quickly varies along pulse longitudes, rapid change of polarization pattern could take place in single pulses.
@@
A natural requirement of our model to work is that none of the orthogonal modes should dominate too largely, in other words, $E_{\mathrm{O}}/E_{\mathrm{E}}$ should be close to 1. Finally, we’d like to emphasize that for explaining the observed polarization rotations, one plasma layer, for which the normal wave modes are elliptical, is equivalent to the two plasma layers described above.
@@ 4.2 Calculations on the phase lag between wave modes
We argue that the phase lag between wave modes plays a key role in the physics behind the observed rapid rotating single pulse polarization orientations. Discussions on relations between polarization orientations and phase lags have been made in Jones (2016, 2020), though they intended to avoid too quick changes in polarization. Observed single pulses’ PA can monotonically rotate by values larger than $\pi$, which requires the phase lags’ difference, between the end longitude and the start longitude, to be larger than $2\pi$, and thus phase lags themselves should be at least of the same order of $2\pi$. This requirement may put some constraints on the magnetospheric parameters.
Theoretical estimations of phase lag start from the dispersion relation of normal modes. We adopt the dispersion relation of O and E modes in a relativistic cold plasma with a single bulk velocity and infinitely large magnetic field Melrose & Stoneham (1977); Arons & Barnard (1986), which could be represented as:
@@
where $n_{\mathrm{O}}$ and $n_{\mathrm{E}}$ are refraction indices of O and E modes, $\theta$ is the angle between wave propagation direction and local magnetic field, and $\omega_{\mathrm{p}}=\sqrt{e^{2}n_{e}/\epsilon_{0}m_{e}}$ is the plasma (angular) frequency. We consider electron-positron plasma here. $\gamma=\sqrt{1/(1-v^{2}/c^{2})}=\sqrt{1/(1-\beta^{2})}$ is the Lorentz factor of magnetospheric particles. Propagation in a trajectory $L$ leads to a phase lag:
@@
$l$ is related to $\theta$, so $n_{\mathrm{O}}$ is the only variable depending on $l$ in the integrand in Eq. 2, according to Eq. 1. In a realistic magnetosphere, $\theta$ changes during propagation because magnetic field’s direction is not fixed along the ray trajectory. In other words, $\theta$ also depends on $l$. Considering the fact that observed rapid rotating polarization orientations in single pulses have very weak dependence on frequency, $n_{\mathrm{O}}$’s frequency dependence, which is actually affected by $\theta$ value, is to be investigated below. Numerical solution of Eq. 1 shown in Figure 6 (ii) tells us that the frequency dependence of $n_{\mathrm{O}}$ (or $\omega(n_{\mathrm{E}}-n_{\mathrm{O}})$) is weakest at $\theta=0^{\circ}$ and strongest at $\theta=90^{\circ}$. The case for very small $\theta$ works better for explaining the weak frequency dependence of rapid polarization rotations we observe.
Given the magnetic field configuration and the distribution of plasma particles, the phase lag between O and E waves of every ray trajectory could be derived. In a dipolar magnetic field, $\theta$ always becomes larger when waves propagate outward, so a small $\theta$ in the whole propagational trajectory prefers the emission to originate relatively close to the magnetic axis, which is also in accord with a small $\beta$ value we’ve mentioned in Section 3.2. For now, we only consider aligned rotator ($\alpha=0^{\circ}$). For oblique rotators, discussions are made in Section 5.1.
Our main assumptions and approximations for the calculation are: (a) pure dipolar magnetic field; (b) ray trajectories are straight lines, where refraction is neglected; (c) wave propagation direction follows only the tangential direction of magnetic field at the emission point; (d) the plasma particle number density decreases as the distance to pulsar center $r$ increases, in the form (similar to Eq. (51) in Ruderman & Sutherland (1975)):
@@
where $n_{\mathrm{GJ,surf}}$ is the Goldreich-Julian number density at pulsar surface (Eq. (9) in Goldreich & Julian (1969)), $B_{\mathrm{surf}}$ is the surface magnetic field strength estimated with period $P$ and period derivative $\dot{P}$ (see Philippov & Kramer, 2022 for example). The neutron star radius $R_{\mathrm{NS}}$ is assumed simply to be 10km.
Integration of equation (2) begins at the emission point and ends at the light cylinder. Given $l$, $n_{\mathrm{O}}$ is determined through calculating $\theta$—then the whole integrand is determined. A schematic diagram of the integration path is shown in Figure 6 (i). With $\kappa=100$, $\gamma=1000$, the calculated phase lags on different frequencies are plotted in Figure 6 (iii). Under these parameter values, $\eta$s lie in the same orders of $2\pi\sim 6\pi$. A larger $\kappa$ or a smaller $\gamma$ lead the phase lag to be larger. When emission height is lower, the phase lag is also larger, and the frequency dependence is weaker. The absence of spectral evolution of observed polarization rotation suggests a small emission height for B1919$+$21.
Figure 6: (i) A schematic diagram of the phase lag calculation. (ii) Results of numerically solving Eq. 1, presented in $\omega\cdot(n_{\mathrm{E}}-n_{\mathrm{O}})$’s dependence on frequency, given different $\theta$. For each curve, the values are normalized with the maximum of the curve, namely the $\omega\cdot(n_{\mathrm{E}}-n_{\mathrm{O}})$ value at 1 GHz. $\gamma$ is chosen $100$, and $n_{e}=10^{9}\mathrm{cm}^{-3}$. (iii) Phase lags (original values and normalized values) between O and E modes versus frequency, with parameters $\kappa=10^{2}$, $\gamma=10^{3}$, $\alpha=0^{\circ}$ and $\phi_{h}=1^{\circ}$. Red, green and blue lines represent results calculated with $h=$ 10, 15 and 20 times of $R_{\mathrm{NS}}$. For details of the theoretical constructions please refer to the text in Section 4.2.

### v2

@@ 3.4 Rapid rotating polarization orientations in single pulses
Next, we directly investigate the individual pulses. Several pulses are presented in Figure 4, with polarization evolution tracks on Poincaré spheres plotted alongside. The polarization states of individual pulses still seem irregular, with PA tracks appearing arbitrary. Not all pulses show “discontinuities” around all three vertical gray dashed lines. Orthogonal mode jumps emerge in some pulses’ core longitude ranges (e.g., #274 in Figure 4, at longitude $\sim 5^{\circ}$). It is worth noting that some pulses present monotonically rotating long PA curves in some longitude ranges (see #7, #275, #555 and #2565 in Figure 4 ,for example). The PA in those pulses rotates by over $180^{\circ}$ or even $360^{\circ}$ across longitudes, accompanied by manifest oscillations of circular polarization (or ellipticity angle) values around 00. The polarization states at these pulse longitudes seem to follow long circular tracks on Poincaré spheres.
Figure 4: Six single pulses with their polarization states are shown on Poincaré spheres. In the “Intensity” panels, the meanings of the lines are the same as those in Figure 1. “$P/I$” panels show the polarization fraction $\sqrt{Q^{2}+U^{2}+V^{2}}/I$ with errorbars on each longitudes (only longitudes where $\sigma_{\mathrm{PA}}\leq 5$ are chosen, same for the “$\psi$” panels and “$\chi$” panels). “$\psi$” panels show PAs with error bars, and “$\chi$” panels show EAs with error bars, on each chosen longitudes. PA and EA dots are also plotted on Poincaré spheres under Hammer-Aitoff projection. For PA dots, EA dots, and the track on Poincaré sphere, the change in color from darker to lighter represents the increase in longitude.
Such rapid rotating polarization patterns are not rare for B1919$+$21, with more than one-third of pulses showing quasi-monotonic rapid rotations of PA in some longitude ranges. This phenomenon seems to depend only very weakly on frequency (see Figure 5 for three examples), at least in our observational bandwidth. In addition, considering PA curves as functions of pulse longitudes, most quasi-monotonic PA curves in single pulses have a generally negative slope (e.g., #7, #555 and #2565 in Figure 4), while only very few of them have a positive slope (e.g., #275 in Figure 4). This asymmetry in PA curve slopes’ distribution might have some physical implications, which we will discuss later. Finally, we would like to point out that the PA curves between two adjacent pulses (e.g., #274 and #275 in Figure 4) can change vastly.
Figure 5: Three single pulses with PA ($\psi$), EA ($\chi$) and total intensity $I$ plotted in two frequency subbands (1000–1125 MHz in green and 1375–1500 MHz in orange). In the $I$ panel, the intensity curves are normalized to the maximum values separately for two subbands. The meanings of the lines and dots are the same as those in Figure 4.
@@ 4 Attempts at modeling
Among the observed phenomena, the rapid rotating polarization orientations in single pulses are outstanding. The PA curves are far from the RVM types. Non-RVM-type PA curves have been discovered and discussed in pulsars (see, e.g., Oswald et al., 2023), magnetars (Lower et al., 2024b), and fast radio bursts (Bera et al., 2024). But the phenomenon in B1919$+$21 is unique in its extreme: in some pulses the PA rotates by over 540 (see #2565 in Figure 4 for example). OPM jumps in integrated profile and single pulses lead us to seek solutions from OPM-related mechanisms. Coherent or partially coherent summation of orthogonally polarized waves (e.g., Dyks, 2019; Dyks et al., 2021; Oswald et al., 2023) is able to reproduce PA curves that are vastly deviated from RVM curves. Our model is an extension of the partial-coherence model by Oswald et al. (2023). In the wave summation models, both changing the wave mode amplitude ratio (mode ratio driven) and changing the phase lag between wave modes (phase lag driven) could result in the polarization vector’s rotation. We will firstly show below that the change of phase lag on different pulse longitudes could be the major reason for the polarization rotation in B1919$+$21’s single pulses. Some supplementary discussions on alternative modeling and model comparisons will be presented in Section 5.1.
@@ 4.1 A phenomenological model: phase lag driven
To reproduce the observed polarization variations in single pulses, here we put forward a phenomenological model based on propagational effects. Based on theories like those in Cheng & Ruderman (1979); Lyubarskii & Petrova (1998); Petrova & Lyubarskii (2000), a pulsar magnetosphere could be divided into the adiabatic walking region (the magnetic field is strong enough and normal modes are linearly polarized) and the polarization limiting region (the weakening or deflection of local magnetic fields introduces more circular polarization components). Although there is no strict boundary between the two regions, we simplify the magnetosphere into two plasma layers, where two initially mutually orthogonal radio waves (O wave and X wave) propagate. Directions of magnetic fields in two layers are fixed separately: in the first layer, the magnetic field is aligned with the O wave electric field vector; in the second layer, the magnetic field has an angle $\theta_{0}$ deflected from the magnetic field of the first layer. The phase lag between the O wave and X wave increases by a value $\eta$ through propagating in the first layer. In the second layer, the original O and X waves are decomposed into new normal wave modes, between which the phase lag increases by $\delta$ through propagating. To include depolarization, we also follow Oswald et al. (2023) to introduce coherence $C$ into our modeling, and then a part of the waves is added incoherently. The mathematical details of our model construction are in Appendix A. Schematic diagrams of the model described above are shown in Figure 6 (i).
@@
Figure 6: (i) A schematic diagram for the phenomenological model. The gray arrows represent magnetic fields. The long green arrow with the label “$\boldsymbol{k}$” means the direction of wave propagation, where refraction is neglected. The blue rectangle with the label “O” and the red parallelogram with the label “X” represent O-mode wave and X-mode wave. (ii) A cross-sectional view of (i), and also a schematic diagram for the difference between two plasma layers. (iii) linear, circular, total polarization fraction, as well as PA and EA on each bin of pulse #555 in Figure 4. (iv) Comparison between pulse #555 (pale dots with error bars) and the pulse simulated through the phenomenological model, with parameters derived from fitting the model to pulse # 555. The details are in the text. The horizontal axis is the number of sample bins, which is equivalent to the longitude. The red dashed line, the blue line, and the cyan line are linear, circular, and total polarization fractions on each bin. The green and magenta dots are PA and EA on each bin.
There are five parameters in the model: (1) initial amplitude ratio $E_{\mathrm{O}}/E_{\mathrm{X}}$; (2) phase lag produced in propagating through the first layer $\eta$; (3) magnetic field deflection angle $\theta_{0}$ (in the second layer, also shown in Figure 6 (i)); (4) phase lag produced in propagating through the second layer $\delta$; and (5) coherence $C$. To fit a pulse with this model, we set $E_{\mathrm{O}}/E_{\mathrm{X}}$, $\theta_{0}$, $\delta$, and $C$ to constants along the pulse longitude, and set $\eta$ to a quadratic function of pulse longitude (or equally, of bin number, $\eta=a_{\eta}(\mathrm{bin}/\mathrm{bin}_{\mathrm{tot}})^{2}+b_{\eta}(\mathrm{bin}/\mathrm{bin}_{\mathrm{tot}})+c_{\eta}$). We apply the modeling to #555 for example. Since our modeling does not care about absolute intensity $I$, we redraw #555 in Figure 6 (iii) with $V/I$, $L/I$ and $P/I$. The first 20 bins, where PA is almost horizontal, are not included in our fitting because they look like one pure mode. We fit $(Q/I,U/I,V/I)$ of #555. The least chi-square fitting result is shown in Figure 6 (iv). Our model could reproduce the main characteristics of the pulse, but performs badly in fitting details. Appropriate functions of $C$ and $E_{\mathrm{O}}/E_{\mathrm{X}}$ versus bin number should improve the fitting, but since the two functions are lack of physical justification at present, we will not go into detail. Anyway, the fitted parameters are: (1) $E_{\mathrm{O}}/E_{\mathrm{X}}=1.01\pm 0.01$; (2) $a_{\eta}=-4.1\pm 0.6$, $b_{\eta}=22.4\pm 0.6$, $c_{\eta}=-2.6\pm 0.2$; (3) $\theta_{0}=77^{\circ}\pm 2^{\circ}$; (4) $\delta=1.2\pm 0.2$; (5) $C=0.390\pm 0.006$. The result indicates that when the phase lag between normal modes rapidly varies along pulse longitudes, rapid change of polarization pattern could take place in single pulses.
@@
A natural requirement for our model to work is that none of the orthogonal modes should dominate too largely, in other words, $E_{\mathrm{O}}/E_{\mathrm{X}}$ should be close to 1 (see Figure 8 for some examples). Finally, we would like to emphasize that to explain the observed polarization rotations, one plasma layer, for which the normal wave modes are elliptical, is equivalent to the two plasma layers described above. The “one plasma layer” version of the propagational model is just the partial-coherence model by Oswald et al. (2023). We introduce two plasma layers for representing different medium properties in different parts of the pulsar magnetosphere, just as we have justified at the beginning of this section.
@@ 4.2 Calculations on the phase lag between wave modes
After all, the magnetosphere is much simplified in either one-plasma or two-plasma models, but we argue that the phase lag between wave modes plays a key role in the physics behind the observed rapid rotating single pulse polarization orientations, which needs further quantitative explorations.
Observed single pulses’ PA can monotonically rotate by values larger than $\pi$, which requires the phase lag difference, between the end longitude and the start longitude, to be greater than $2\pi$, and thus phase lags themselves should be at least on the same order of $2\pi$. This requirement may put some constraints on the magnetospheric parameters.
Theoretical estimations of phase lag start from the dispersion relation of normal modes. We adopt the dispersion relation of O and X modes in a relativistic cold plasma with a single bulk velocity and infinitely large magnetic field (Melrose & Stoneham, 1977, Arons & Barnard, 1986), which could be represented as:
@@
where $n_{\mathrm{O}}$ and $n_{\mathrm{X}}$ are the refraction indices of O and X modes, $\theta$ is the angle between wave propagation direction and local magnetic field, and $\omega_{\mathrm{p}}=\sqrt{e^{2}n_{e}/\epsilon_{0}m_{e}}$ is the plasma (angular) frequency. We consider the electron-positron plasma here. $\gamma=\sqrt{1/(1-v^{2}/c^{2})}=\sqrt{1/(1-\beta^{2})}$ is the Lorentz factor of magnetospheric particles. Propagation on a trajectory $L$ leads to a phase lag:
@@
where $k_{\mathrm{X}}$ and $k_{\mathrm{O}}$ refer to the wavenumbers of the X and O waves. $l$ is related to $\theta$, so $n_{\mathrm{O}}$ is the only variable depending on $l$ in the integrand in Equation 9, according to Equation 8. In a realistic magnetosphere, $\theta$ changes during propagation because the magnetic field’s direction is not fixed along the ray trajectory. In other words, $\theta$ also depends on $l$. Considering the fact that observed rapid rotating polarization orientations in single pulses have a very weak dependence on frequency, $n_{\mathrm{O}}$’s frequency dependence, which is actually affected by the $\theta$ value, is to be investigated below. The numerical solution of Equation 8 shown in Figure 7 (ii) tells us that the frequency dependence of $n_{\mathrm{O}}$ (or $\omega(n_{\mathrm{X}}-n_{\mathrm{O}})$) is weakest at $\theta=0^{\circ}$ and strongest at $\theta=90^{\circ}$. The case for very small $\theta$ works better for explaining the weak frequency dependence of rapid polarization rotations that we observe.
Given the magnetic field configuration and the distribution of plasma particles, the phase lag between O and X waves of every ray trajectory could be derived. In a dipolar magnetic field, $\theta$ always becomes larger when waves propagate outward, so a small $\theta$ in the whole propagational trajectory prefers the emission to originate relatively close to the magnetic axis, which is also in accord with the small $\beta$ value we have mentioned in Section 3.2. For now, we only consider the aligned rotator ($\alpha=0^{\circ}$). For oblique rotators, discussions are made in Section 5.1.
Our main assumptions and approximations for the calculation are (a) pure dipolar magnetic field; (b) ray trajectories are straight lines, where refraction is neglected; (c) wave propagation direction follows only the tangential direction of magnetic field at the emission point; and (d) the plasma particle number density decreases as the distance to pulsar center $r$ increases, in the form (similar to Equation (51) in Ruderman & Sutherland (1975)):
@@
where $n_{\mathrm{GJ,surf}}$ is the Goldreich-Julian number density at pulsar surface (Equation (9) in Goldreich & Julian (1969)), $B_{\mathrm{surf}}$ is the surface magnetic field strength estimated with period $P$ and period derivative $\dot{P}$ (see Philippov & Kramer, 2022 for example). The neutron star radius $R_{\mathrm{NS}}$ is assumed simply to be 10km.
The integration of Equation 9 begins at the emission point and ends at the light cylinder. Given $l$, $n_{\mathrm{O}}$ is determined through calculating $\theta$—then the whole integrand is determined. A schematic diagram of the integration path is shown in Figure 7 (i). With $\kappa=100$, $\gamma=1000$, the phase lags calculated at different frequencies are plotted in Figure 7 (iii). Under these parameter values, $\eta$s lie on the same orders of $2\pi\sim 6\pi$. A larger $\kappa$ or a smaller $\gamma$ cause the phase lag to be larger. When emission height is lower, the phase lag is also larger, and the frequency dependence is weaker. The absence of spectral evolution of observed polarization rotations suggests a small emission height for B1919$+$21.
Figure 7: (i) A schematic diagram of the phase lag calculation. (ii) Results of numerically solving Equation 8, presented in $\omega\cdot(n_{\mathrm{X}}-n_{\mathrm{O}})$’s dependence on frequency, given different $\theta$. For each curve, the values are normalized with the maximum of the curve, namely, the $\omega\cdot(n_{\mathrm{X}}-n_{\mathrm{O}})$ value at 1 GHz. $\gamma$ is chosen to be $100$, and $n_{e}=10^{9}\mathrm{cm}^{-3}$. (iii) Phase lags (original values and normalized values) between O and E modes versus frequency, with parameters $\kappa=10^{2}$, $\gamma=10^{3}$, $\alpha=0^{\circ}$ and $\phi_{h}=1^{\circ}$. Red, green, and blue lines represent results calculated with $h=$ 10, 15, and 20 times $R_{\mathrm{NS}}$. For details of the theoretical constructions, please refer to the text in Section 4.2.